OpenStudy (anonymous):

Can someone help me to rationalize the following expression, please? 3/(1-sqrt(12))

OpenStudy (lalaly):

\[\frac{3}{1-\sqrt{12}} \times \frac{1+\sqrt{12}}{1+\sqrt{12}}= \frac{3+3\sqrt{12}}{1-12}\]

OpenStudy (lalaly):

\[=\frac{3+3\sqrt{12}}{-11}\]

OpenStudy (anonymous):

Thanks

OpenStudy (lalaly):

=)